OpenStudy (kittykatiehunt):

The function g(x) = 3x2 − 12x + 7 written in vertex form is g(x) = 3(x − 2)2 − 5. What is the vertex of g(x)? (−6, −5) (−2, −5) (2, −5) (6, −5)

OpenStudy (google):

you still need help?

OpenStudy (trojanpoem):

y = a(x - h)^2 + k by comparing coefficients: h= 2 , k = -5 so vertex = (h k) = (2, -5)